The sum of the ages of Dada and his brother is 22. Five years ago, Dada was 3 times as old as his brother.
How old are they now?

Answer by ikleyn(52781)

.

So, 5 years ago the sum of their ages was  22-5-5 = 12, and one was 3 times as old as the other.


Hence, 5 years ago they were  3 and 9 years old;


It means that now they are 3+5 = 8 and 9+5 = 14 years old.    ANSWER

Mental solution without using equations.

------------------

There is a bunch of lessons on age word problems in this site:
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - One unusual age word problem
    - Miscellaneous age problems
    - Age problem for the day of April, 1
    - OVERVIEW of lessons on age problems

Read them attentively and become an expert in this field.